# SciLitLLM: How to Adapt LLMs for Scientific Literature Understanding

### v2

@@ 3.1.1. Format & Grammar Correction
As illustrated in Appendix A, a parsed text from a PDF document often contains many formatting and syntax errors. To address this issue, we prompt a moderate yet powerful language model, Llama3-7B-Instruct (AI@Meta, 2024), to correct these errors introduced during the PDF parsing process. Utilizing the vLLM (Kwon et al., 2023) backend, Llama3-7B-Instruct can process approximately 2.52 million tokens per Nvidia A100 GPU hour. The process takes over 5,000 A100 GPU hours to handle all 73,000 textbooks and 625,000 research papers. Example texts – both before and after processing – along with the prompt template are provided in Appendix A to demonstrate the improvements made through this correction process.
3.1.2. CPT Quality Filter
@@
During CPT, maintaining the quality of the training corpus is crucial for effective knowledge injection. Given the extremely large scale of pre-training corpora, assessing quality through human annotation is not feasible (Erdmann et al., 2019; Qiu et al., 2023). Consequently, leading LLMs (e.g., Phi (Gunasekar et al., 2023), Llama (Touvron et al., 2023), and Qwen (Yang et al., 2024)) employ model-based quality filters. The typical process involves using larger LLMs to score the quality of a subset of texts, which then serve as labels for training small classifiers (e.g., random forest (Gunasekar et al., 2023) and Bert (AI@Meta, 2024)) to annotate the entire training corpus. Inspired by this approach, we design a resource-efficient method based on a lightweight text quality classifier.
@@
Following prior studies (Gunasekar et al., 2023; AI@Meta, 2024; Yang et al., 2024; Anton et al., 2024), we first annotate a random subset of 50k CPT texts using a powerful model – Llama3-70B-Instruct (AI@Meta, 2024). We adapt the quality assessment prompt from fineweb-edu-classifier (Anton et al., 2024), a widely-used quality classifier for web data, to evaluate the educational value (Gunasekar et al., 2023) of the scientific knowledge in each sampled text, assigning scores ranging from 0 (lowest quality) to 5 (highest quality). After annotation, we perform supervised transfer learning on the fineweb-edu-classifier (Anton et al., 2024) checkpoint – a Bert-based (Devlin et al., 2019) quality classifier in the web domain. This process results in a scientific text quality classifier tailored for scientific corpus assessment. See Appendix B for more details about classifier training and hyperparameters.
We then utilize this classifier to assess the quality of the entire CPT dataset (See Figure 3 for concrete samples). Each sample is evaluated and assigned with a continuous real number as the quality score. To enhance the overall quality of training data, we then exclude the lowest-scoring 25% from the dataset.
By leveraging the CPT quality classifier, we can efficiently filter out low-quality texts and ensure that only high-quality, informative content is retained. This step is crucial for enhancing the scientific knowledge base of our LLM, thereby improving its performance in scientific literature understanding.
3.1.3. CPT Training Settings
We perform CPT on Qwen2-Base (Yang et al., 2024) for one epoch, encompassing 23.7 billion tokens (cf. Table 1), with a sequence length of 2,048 tokens. To maintain the model’s general knowledge, we also include a similar scale of general corpus tokens from Redpajama (Computer, 2023). To stabilize the learning procedure, we gradually decrease the learning rate from $1\times 10^{-5}$ to 0 for SciLitLLM-7B, and from $5\times 10^{-6}$ to 0 for SciLitLLM-72B (QLoRA), with a cosine scheduler. To address overfitting, we apply a weight decay of $0.1$ and gradients were clipped at a maximum value of $1.0$. The CPT training took approximately 3 days on 32 Nvidia A100 GPUs for SciLitLLM-7B-Base (full parameters) and about 10 days for SciLitLLM-72B-Base (QLoRA).
@@ 3.1.3. CPT Training Settings
Infinity-Instruct [^5] https://huggingface.co/datasets/BAAI/Infinity-Instruct General 3M 1.7B
Table 1. Data statistics of continual pre-training and supervised fine-tuning. #Doc/#Ins denotes the number of documents of CPT corpora and the number of instructions for SFT, respectively. Underlined datasets are curated by us.
3.2. SFT for Scientific Instruction Following
@@
  • Existing instruction-tuning datasets in the scientific domain (Feng et al., 2024; Fang et al., 2023; Li et al., 2024) primarily focus on fields such as physics, chemistry, and biology. Manually collecting instruction-tuning data for other less-represented vertical domains (e.g., alloy, biomedicine, and material) is both time-consuming and costly (Erdmann et al., 2019; Qiu et al., 2023).
@@
3.2.1. Instruction Synthesis of Less-represented Domains
Unlike typical question-answer pairs, an instruction for a scientific literature understanding task comprises three components: (1) a segment of scientific literature, (2) a question pertaining to the context, and (3) the corresponding answer (Wadden et al., 2024). Simply prompting an LLM to generate a scientific context along with an associated question-answer pair – without variations in the instructions or parameters – often yields similar or repeated contents. This phenomenon arises because language models tend to adhere to the most probable or common paths dictated by their memory base and priors, thereby lacking the creativity to explore diverse generation (Gunasekar et al., 2023). Consequently, we are motivated to devise a strategy to encourage the language model to produce more creative and diverse instructions for scientific literature understanding while simultaneously ensuring the quality and coherence of the generated contexts.
We design a simple yet effective three-step pipeline to generate diverse and high-quality instructions for scientific contexts and corresponding question-answer pairs, consisting of the following:
  1. (1)
    Probability table of domain keywords. For a target scientific domain (e.g., alloy, biomedicine, and material), we collect dozens of high-impact research papers via Google Scholar [^6] https://scholar.google.com/ and count the frequency of each word appearing in these papers. Subsequently, we remove spaces and meaningless articles such as “a,” “an,” “the,” etc. After normalization, a probability table of domain keywords, representing the word-level distribution of domain literature, is obtained.
  2. (2)
    Scientific task descriptions. Since LLMs are expected to handle various types of scientific tasks, an instruction set with task diversity is essential. Therefore, we compile a list of task descriptions by including representative tasks from existing scientific NLP datasets (Wadden et al., 2024; Cai et al., 2024; Feng et al., 2024), covering as many scenarios as possible that an LLM may encounter in real applications.
  3. (3)
@@ 3.2.1. Instruction Synthesis of Less-represented Domains
The detailed generation process and example prompts are presented in Appendix C.1. Utilizing this pipeline, we obtain over 100k synthetic instructions for scientific literature understanding, covering less-represented scientific domains and various types of specialized tasks.
3.2.2. Instruction Quality Control
Figure 5. The quality of SciLitInsis evaluated from five aspects: clarity, complexity, correctness, usefulness, and adaptability (the higher the better). Instructions with an average score of less than 4 are filtered out.
@@ 3.2.2. Instruction Quality Control
    Heuristic deduplication. Despite the measures taken during the generation process to prevent high homogeneity in the instructions, the generated data points may still contain similar questions or identical answers. To eliminate such redundancy, we implement a simple yet effective deduplication process using the Levenshtein distance to calculate the similarity score between instructions. Based on this score, $5\%$ to $10\%$ of similar data are removed for each question type. Detailed processing steps are provided in Appendix C.2.
  2. (2)
@@
    LLM-based filtering. Inspired by recent efforts (Eldan and Li, 2023; Chiang and Lee, 2023; Zhang et al., 2024c) to measure the quality of generated content using LLMs, we leverage Llama-3-70B-Instruct (AI@Meta, 2024) to assess the quality of generated instructions for scientific literature understanding. Specifically, the quality is evaluated from five aspects: clarity, complexity, correctness, usefulness, and adaptability, assigning each instruction a score from 0 (lowest quality) to 5 (highest quality). Instructions with an average score of less than 4 are filtered out. We show the quality statistics of synthetic instructions in Figure 5. The detailed recipe for instruction quality evaluation with concrete examples is included in Appendix C.3.
@@
Through instruction synthesis and quality control pipeline, we obtain SciLitIns, consisting of approximately 110,000 high-quality and diverse instructions for scientific literature understanding. With these instructions, models’ problem-solving abilities in this specialized field could be enhanced.
3.2.3. SFT Training Settings
@@
Our SFT training dataset consists of three parts: SciLitIns, SciRIFF (Wadden et al., 2024) and Infinity-Instruct [^7] https://huggingface.co/datasets/BAAI/Infinity-Instruct, as shown in Table 1. Infinity-Instruct is a collection of more than twenty open-source instructions datasets, covering various general domains. SciRIFF and SciLitIns contain specialized instructions for scientific literature understanding. We use full parameter training for SciLitLLM-7B-Base and QLoRA (Dettmers et al., 2023) parameter-efficient training for SciLitLLM-72B-Base. For both models, we train for one epoch on Infinity-Instruct to cultivate their general instruction-following abilities, then for five epochs on SciLitIns and SciRIFF for scientific literature understanding enhancement. The training is conducted with a sequence length of 4,096, a maximum learning rate of $5\times 10^{-6}$, and a cosine scheduler. The SFT training takes approximately 32 hours on 32 A100 GPUs for the 7B and 100 hours for the 72B model, resulting in SciLitLLM-7B-Instruct and SciLitLLM-72B-Instruct.
Models MMLU-s CMMLU-s Xiezhi-en-s Xiezhi-ch-s
# Parameters ¡ 15B
ChatGLM-6B 46.99 46.42 58.33 63.50
Mistral-7B 59.84 36.43 64.97 53.34
Qwen1.5-7B 57.09 73.74 65.87 71.60
Qwen2-7B 66.62 86.04 71.53 74.36
Llama2-7B 39.87 28.33 42.40 36.40
Llama3-8B 61.52 42.16 66.29 63.46
Llama2-13B 49.06 32.29 58.30 42.65
Qwen1.5-14B 66.67 80.19 68.60 73.94
SciLitLLM-7B 70.85 91.84 73.42 78.24
# Parameters ¿ 50B
Mixtral-8x7B 67.58 44.88 69.54 65.81
Qwen2-57B-A14B 73.79 89.65 70.35 73.73
Llama2-70B 65.03 43.94 66.75 65.98
Llama3-70B 76.43 66.41 71.36 73.28
Qwen1.5-72B 74.89 86.87 71.18 74.47
Qwen2-72B 80.86 92.31 72.41 75.03
Qwen1.5-110B 77.06 90.72 73.45 73.14
SciLitLLM-72B (QLoRA) 82.31 93.08 74.23 76.93
Table 2. Performance comparison of base models. Bold indicates the highest performance for LLMs under 15B parameters or above 50B parameters. SciLitLLM achieves leading performance on all four scientific knowledge benchmarks.
Dataset Domain/ Task # Parameter ~7B # Parameter ~70B API
SciTulu-7B Mistral-7B Llama3-8B Qwen2-7B SciLitLLM-7B Llama3-70B Qwen2-72B SciLitLLM-72B (QLoRA) GPT3.5 GPT4o
SciAssess FundSci 32.3 48.3 58.5 70.3 74.8 70.9 77.1 78.4 62.2 76.7
AlloyMat 23.9 28.0 32.9 32.8 35.6 44.9 42.7 49.1 32.0 52.1
Biomed 67.8 76.0 77.4 80.8 79.6 79.6 81.0 79.6 78.0 82.3
DrugDisc 25.4 30.2 32.0 31.7 33.2 41.5 35.5 41.8 31.0 43.4
OrgMat 16.7 20.6 24.5 28.3 38.9 41.5 52.7 48.6 24.4 62.7
@@
SciRIFF BioASQ 37.5 43.9 44.7 40.7 51.0 46.3 43.6 50.7 47.3 46.7
BioR 55.7 48.2 45.3 44.3 74.0 59.9 59.1 63.0 53.9 61.0
DiscMT 61.5 44.6 58.7 59.9 77.4 71.9 73.5 73.5 67.9 78.3
EI 11.6 17.1 14.7 14.4 22.3 22.0 24.1 23.1 19.2 24.7
MC 34.6 47.0 49.5 51.6 68.0 59.9 57.3 70.5 47.8 58.7
MuP 72.1 93.4 90.7 96.6 76.8 96.4 97.8 77.9 76.8 86.9
Qasper 54.2/38.6 58.6/39.4 58.2/41.9 56.8/34.5 58.4/56.9 25.0/19.4 63.3/47.2 60.5/54.1 54.7/39.8 67.8/50.5
SciERC 35.6 30.2 19.9 27.5 39.9 35.2 34.1 46.9 28.6 42.2
SciFact 66.0/49.2 68.5/51.3 64.6/51.7 65.2/44.3 68.5/59.7 85.1/67.3 82.3/65.9 75.2/60.6 69.7/53.3 84.3/68.7
@@
Table 3. Model performances on scientific literature understanding benchmarks: SciAssess and SciRIFF. SciLitLLM-7B and SciLitLLM-72B achieve leading performance compared with models of similar scales. The best-performing models in the 7B and 70B scales are highlighted in bold. Results for SciTulu-7B, GPT-3.5, and GPT-4o on SciRIFF are taken from its original papers, while all other results are generated by our experiments.
4. Experiments
In this section, we perform experiments to answer the following research questions:
  • RQ1: How does SciLitLLM perform on scientific literature understanding tasks?
  • RQ2: Can CPT with domain-specific corpora aid in scientific knowledge injection?
  • RQ3: Can SFT with synthetic instructions improve performance on scientific literature understanding tasks?
4.1. Experimental Setup
4.1.1. Benchmarks
@@
  • CPT benchmarks. We evaluate the base models on three widely adopted benchmarks: MMLU (Hendrycks et al., 2021), CMMLU (Li et al., 2023b), and Xiezhi (Gu et al., 2024). Specifically, we select the STEM subsets from these benchmarks to assess their scientific knowledge, which serves as the foundation for scientific literature understanding.
@@
  • SFT benchmarks. We evaluate the instruct models on scientific literature understanding benchmarks: SciRIFF (Wadden et al., 2024) and SciAssess (Cai et al., 2024). Brief descriptions of them are provided in Appendix D.
4.1.2. Baselines
We test the following baselines:

### v3

@@ 3.1.1 Format & Grammar Correction
A parsed text from a PDF document often contains many formatting and syntax errors. To address this issue, we prompt Llama3-8B-Instruct, to correct these errors introduced during the PDF parsing process. Utilizing the vLLM (Kwon et al., 2023) backend, Llama3-8B-Instruct can process approximately 2.52 million tokens per Nvidia A100 GPU hour. The process takes over 5,000 A100 GPU hours to handle all the textbooks and research papers. Example texts – both before and after processing – along with the prompt template are provided in Appendix A to demonstrate the improvements through this correction process.
3.1.2 CPT Quality Filter
@@
During CPT, maintaining the quality of the training corpus is crucial for effective knowledge injection. Given the extremely large scale of pre-training corpora, assessing quality through human annotation is not feasible (Erdmann et al., 2019; Qiu et al., 2023). Consequently, leading LLMs (e.g., Phi (Gunasekar et al., 2023), Llama (Touvron et al., 2023), and Qwen2.5 (Team, 2024)) employ model-based quality filters. The typical process involves using larger LLMs to score the quality of a subset of texts, which then serve as labels for training small classifiers (e.g., random forest (Gunasekar et al., 2023) and Bert (AI@Meta, 2024)) to annotate the entire training corpus. Inspired by this approach, we design a resource-efficient method based on a lightweight text quality classifier.
@@
We first annotate a random subset of 50k CPT texts using Llama3-70B-Instruct. We adapt the quality assessment prompt from fineweb-edu-classifier (Anton et al., 2024), a widely-used quality classifier for web data, to evaluate the educational value (Gunasekar et al., 2023) of the scientific knowledge in each sampled text, assigning scores ranging from 0 (lowest quality) to 5 (highest quality). After annotation, we perform supervised transfer learning on the fineweb-edu-classifier checkpoint.
This process results in a scientific text quality classifier tailored for scientific corpus assessment. We then utilize this classifier to assess the quality of the entire CPT dataset and exclude the lowest-scoring 25% (See Figure 4 for concrete samples). See Appendix B for more details about classifier training and hyperparameters.
3.1.3 CPT Training Settings
@@
Infinity-Instruct General 3M 1.7B
Table 1: Data statistics of CPT and SFT stages. Underlined datasets are curated by us.
After preparing all the corpus, we perform CPT on Qwen2.5-Base (Team, 2024) for one epoch, encompassing 23.7 billion tokens (cf. Table 1), with a sequence length of 2,048 tokens. To maintain the model’s general knowledge, we also include a similar scale of general corpus tokens from Redpajama (Computer, 2023). To stabilize the learning procedure, we gradually decrease the learning rate from $1\times 10^{-5}$ to 0 with a cosine scheduler. To address overfitting, we apply a weight decay of $0.1$ and gradients were clipped at a maximum value of $1.0$. The CPT training took approximately 3 days on 32 Nvidia A100 GPUs for SciLitLLM-7B-Base and about 7 days for the 14B model.
3.2 SFT for Scientific Instruction Following
@@
  • Existing instruction-tuning datasets in the scientific domain (Feng et al., 2024; Fang et al., 2023; Li et al., 2024) primarily focus on fields such as physics, chemistry, and biology. Manually collecting instruction-tuning data for other less-represented vertical domains (e.g., biomedicine, and material) is both time-consuming and costly (Erdmann et al., 2019; Qiu et al., 2023).
@@
3.2.1 Instruction Synthesis of Less-represented Domains
Unlike typical question-answer pairs, an instruction for a scientific literature understanding task comprises three components (Wadden et al., 2024): (1) a segment of scientific literature, (2) a question pertaining to the context, and (3) the corresponding answer. Simply prompting an LLM to generate a scientific context along with an associated question-answer pair – without variations in the instructions or parameters – often yields similar or repeated contents. Thus, we design a simple yet effective three-step pipeline to generate diverse and high-quality scientific contexts and corresponding question-answer pairs, consisting of the following:
  1. 1.
    Probability table of domain keywords. For a target scientific domain (e.g., biomedicine, and material), we collect dozens of high-impact research papers via Google Scholar [^6] https://scholar.google.com/ and count the word frequency appearing in these papers. Then, we obtain a probability table of domain keywords.
  2. 2.
    Scientific task list. Since LLMs are expected to handle various types of scientific tasks, an instruction set with task diversity is essential. Therefore, we compile a list of task descriptions by including representative tasks from existing scientific NLP datasets (Wadden et al., 2024; Cai et al., 2024; Feng et al., 2024), covering as many scenarios as possible that an LLM may encounter in real applications.
  3. 3.
@@ 3.2.1 Instruction Synthesis of Less-represented Domains
Utilizing this pipeline, we obtain over 100k synthetic instructions covering less-represented scientific domains and various types of specialized tasks (details are presented in Appendix C.1).
3.2.2 Instruction Quality Control
@@
    Heuristic deduplication. Despite the measures taken during the generation process to prevent high homogeneity in the instructions, the generated data points may still contain similar questions or identical answers. To eliminate such redundancy, we implement a deduplication process to remove similar data points. See Appendix C.2 for details.
  2. 2.
@@
    LLM-based filtering. Inspired by recent efforts (Eldan & Li, 2023; Chiang & Lee, 2023; Zhang et al., 2024c) to measure the quality of generated content using LLMs, we leverage Llama-3-70B-Instruct to assess the quality of generated instructions for five aspects: clarity, complexity, correctness, usefulness, and adaptability. We show the quality statistics of synthetic instructions in Figure 7. The detailed recipe for instruction quality evaluation with concrete examples is included in Appendix C.3.
@@
Through instruction synthesis and quality control pipeline, we obtain SciLitIns, consisting of 93,894 high-quality and diverse instructions for scientific literature understanding.
3.2.3 SFT Training Settings
@@
Our SFT training dataset consists of three parts: SciLitIns, SciRIFF (Wadden et al., 2024) and Infinity-Instruct [^7] https://huggingface.co/datasets/BAAI/Infinity-Instruct, as shown in Table 1. Infinity-Instruct is a collection of more than twenty open-source instructions datasets, covering various general domains. SciRIFF and SciLitIns contain specialized instructions for scientific literature understanding. For both scales, we train for one epoch on Infinity-Instruct to cultivate their general instruction-following abilities, then for five epochs on SciLitIns and SciRIFF for scientific literature understanding enhancement. The training is conducted with a sequence length of 4,096, a maximum learning rate of $1\times 10^{-5}$, and a cosine scheduler. The SFT training takes approximately 32 hours for the 7B and 70 hours for the 14B model on 32 A100 GPUs, resulting in SciLitLLM-7B-Instruct and SciLitLLM-14B-Instruct.
4 Experiments
Models MMLU-Pro-Bio MMLU-Pro-Chem MMLU-Pro-Heal MaScQA Avg.
# Parameters $<$ 10B
Llama3.1-8B 56.35 25.88 44.50 53.28 45.00
Qwen2.5-7B 68.36 48.37 51.64 56.23 56.15
SciLitLLM-7B 70.45 51.21 54.06 59.27 58.75
# Parameters $>$ 10B
Qwen2.5-14B 79.78 62.73 62.22 66.92 67.91
SciLitLLM-14B 80.62 64.44 64.24 68.06 69.34
Table 2: Performance comparison of base models. SciLitLLM outperforms Llama3.1 and Qwen2.5 with similar scales, demonstrating improved domain-specific understanding through continual pretraining on scientific corpora.
Dataset Domain/ Task # Parameter $<$ 10B # Parameter $>$ 10B API
SciTulu-7B Mistral-7B Llama3.1-8B Qwen2.5-7B SciLitLLM-7B Qwen2.5-14B Llama3.1-70B Qwen2.5-72B SciLitLLM-14B GPT3.5 GPT4o
SciAssess Biology 45.3 52.0 63.4 60.9 65.3 65.9 69.4 67.2 67.0 55.4 68.9
Chemistry 19.0 34.0 46.1 47.9 55.4 60.0 62.5 63.7 63.4 37.3 68.0
Material 31.3 36.5 50.9 48.9 53.7 57.9 59.2 59.4 62.6 37.0 62.0
Medicine 19.9 25.1 30.2 28.3 32.4 31.5 37.4 37.8 39.9 31.8 45.8
@@ 4 Experiments
SciRIFF BioASQ 37.5 43.2 45.0 45.3 51.2 48.0 43.5 46.7 53.9 47.3 46.7
BioR 55.7 48.9 48.1 43.9 76.6 56.8 61.5 57.8 83.3 53.9 61.0
DiscMT 61.5 44.5 68.8 73.7 71.1 70.5 77.3 54.7 77.7 67.9 78.3
EI 11.6 17.2 17.2 23.3 23.5 27.2 26.0 26.3 30.8 19.2 24.7
MC 34.6 47.5 49.7 50.8 70.7 54.2 60.9 58.0 72.1 47.8 58.7
MuP 72.1 87.5 87.0 84.9 67.5 91.5 90.9 94.0 67.7 76.8 86.9
Qasper 54.2/38.6 48.8/43.2 46.5/42.7 38.1/25.7 50.7/54.1 48.4/46.8 49.1/40.2 61.9/50.5 54.0/55.1 54.7/39.8 67.8/50.5
SciERC 35.6 31.1 32.9 30.1 49.9 33.7 36.2 34.1 54.7 28.6 42.2
SciFact 66.0/49.2 68.8/53.8 63.8/53.7 77.8/60.3 83.5/67.3 83.1/65.6 87.3/70.3 86.5/69.6 91.0/68.6 69.7/53.3 84.3/68.7
@@
Table 3: Model performances on scientific literature understanding benchmarks: SciAssess and SciRIFF. SciLitLLM-7B and SciLitLLM-14B achieve leading performance compared with open-source models of similar scales. The best-performing models smaller than 10B and larger than 10B are highlighted in bold. Results for SciTulu-7B, GPT-3.5, and GPT-4o on SciRIFF are taken from its original papers, while all other results are produced by us.
In this section, we perform experiments to answer the following research questions: (Q1) How does SciLitLLM perform on scientific literature understanding tasks? (Q2) Can CPT with domain-specific corpora aid in scientific knowledge injection? (Q3) Can SFT with SciLitIns improve performance on scientific literature understanding tasks?
4.1 Experimental Setup
4.1.1 Benchmarks
@@
  • Base model benchmarks. To assess their scientific knowledge, which serves as the foundation for scientific literature understanding, we evaluate the base models on scientific subjects – biology, chemistry, and health in MMLU-Pro (Wang et al., 2024) and MaScQA (Zaki et al., 2023) – a question answering dataset of materials science.
@@
  • Instruction model benchmarks. We evaluate the instruct models on scientific literature understanding benchmarks: SciRIFF (Wadden et al., 2024) and SciAssess (Cai et al., 2024). Brief descriptions of them are provided in Appendix D.
4.1.2 Baselines